Determine (-2+i3)^5 in the polar form and in cartesian form
Found 2 solutions by stanbon, Edwin McCravy:
Answer by stanbon(75887)

Determine (-2+i3)^5 in the polar form
---
r = sqrt(2^2+3^2) = sqrt(13)
theta = tan^-1(-3/2) = 303.69 degrees
----------------
(-2+3i)^5 = 13^(5/2)[cos(303.69)+isin(303.69)
=====================================================
and in cartesian form
(-2+3i)^5 = x + yi
x = 13^(5/2)*cos(303.69) = 338
y = 13^(5/2)*isin(303.69) = -507i
-----------------------------------------
Cheers,
Stan H.
==========

Answer by Edwin McCravy(20055)

Determine (-2+i3)^5 in the polar form and in cartesian form

The other tutor got the angle in the 4th quadrant instead
of the 2nd quadrant, and also forgot to multiply the angle
by 5, which also would have been wrong since he got the
angle in the wrong quadrant.

Draw the vector x + iy = -2 + i3 which is the 
vector from (0,0) to (-2,3), and let its length 
(magnitude, modulus, absolute value) be r, and its
angle (argument) be q:



Then we draw a perpendicular (in green) from the
tip of the vector to the x-axis:


 
x + iy = eiq = r⋅cis(q) = r⋅(cosq + i⋅sinq)

We calculate r and q)

r² = x² + y²
r² = (-2)² + (3)²
r² = 4 + 9
r² = 13
 r = √13

tanq = y/x
tanq = (3)/(-2) = -1.5
q = 123.69° = 2.16 radians

We use:
(x + iy)n = rnei·nq = rncis(nq) = rn[cos(nq) + i·sin(nq)]

with n = 5, q = 123.69°, r = √13

(√13)5 = 609.3381656

5q = 5(123.69°) = 618.45° which
is coterminal with 258.45°

609.3381656[cos(258.45°) + i·sin(258.45°)] =  -122-i·597
 
Edwin